OpenStudy (markkboyy):

the following is a geometric sequence 3,4,7,9... true or false

MARC (marc_d):

the following sequence is a geometric sequence if the common ratio,r is equal\[T_1=3\]\[T_2=4\]\[T_3=7\]\[T_4=9\]\[common~ratio,r=\frac{ T_2 }{ T_1 }\]\[r=\frac{ 4 }{ 3 }\]\[r=\frac{ T_3 }{ T_2 }\]\[r=\frac{ 7 }{ 4 }\]\[r=\frac{ T_4 }{ T_3 }\]\[r=\frac{ 9 }{ 7 }\]

MARC (marc_d):

we can see that\[\frac{ T_2 }{ T_1 }\color{red}{\neq} \frac{ T_3 }{ T_2 }\color{red}{\neq} \frac{ T_4 }{ T_3 }\]

MARC (marc_d):

If the common ratio are not equal,therefore,it is not a geometric sequence

OpenStudy (markkboyy):

so false

MARC (marc_d):

yes!

OpenStudy (markkboyy):

can you help me on some more of these and ill keep giving more medals?

MARC (marc_d):

i will try my best..tag me..

OpenStudy (markkboyy):

okay